# GenCode: A Generic Data Augmentation Framework for Boosting Deep Learning-Based Code Understanding

### v2

@@ Introduction
Table 6: Effectiveness of data augmentation methods w.r.t. test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on original test data and ASR $\downarrow$ (%) on test data. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
@@
Model DA method Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$
GenCode (Random) 94.74 ± 0.21 25.21 50.32 98.82 ± 0.15 27.65 36.65 98.24 ± 0.11 32.52 44.87 98.27 ± 0.19 8.01 18.12
GenCode (Min) 93.34 ± 0.32 35.66 66.67 96.13 ± 0.21 65.51 69.76 96.43 ± 0.17 46.76 53.86 97.23 ± 0.14 20.65 41.32
CodeBERT GenCode 96.23 ± 0.18 22.33 42.62 99.32 ± 0.24 22.75 29.25 98.89 ± 0.09 29.89 40.65 98.93 ± 0.15 7.12 15.83
GenCode (Random) 95.26 ± 0.11 29.97 50.63 98.93 ± 0.18 17.56 24.29 98.68 ± 0.22 20.65 38.86 98.67 ± 0.13 2.04 5.97
GenCode (Min) 93.53 ± 0.17 40.55 68.23 96.31 ± 0.24 38.73 46.98 96.72 ± 0.12 26.87 46.67 97.22 ± 0.21 11.45 21.23
GraphCodeBERT GenCode 97.93 ± 0.14 26.43 39.87 99.45 ± 0.17 15.78 20.92 99.16 ± 0.13 18.76 33.81 99.23 ± 0.09 1.86 4.62
GenCode (Random) 96.24 ± 0.21 25.03 47.61 98.99 ± 0.21 18.46 23.67 98.93 ± 0.16 25.11 42.34 98.83 ± 0.08 8.67 17.42
GenCode (Min) 94.11 ± 0.31 38.65 59.89 98.29 ± 0.12 35.67 48.59 97.03 ± 0.23 32.61 49.27 97.01 ± 0.15 18.36 25.54
CodeT5 GenCode 98.85 ± 0.16 23.57 38.76 99.64 ± 0.37 12.13 18.81 99.38 ± 0.11 21.63 39.76 99.34 ± 0.12 7.39 12.15
Visualizing principal component analysis. Furthermore, we explore why models trained by GenCode have higher accuracy than models trained using other methods by analyzing the code representation capability of trained code embeddings. Specifically, 1) first, we extract the code embeddings produced by each trained model using all the test data as code features. 2) Then, since the extracted features are high dimensions and hard to analyze, we reduce the dimensions of code data features using the Principal Component Analysis (PCA) (Maćkiewicz and Ratajczak 1993) algorithm. 3) After that, we visualize the reduced 2-dimensional vectors. Normally, the features extracted by a well-trained model are more clearly distributed based on the data labels. Figure 3 depicts the visualized features extracted from CodeBERT using the Refactory dataset. We can see that the code embeddings trained by using GenCode are more clearly distributed than embedding extracted by other models, i.e., the buggy code and correct code are greatly different in the figure. These results indicate that GenCode can train good code embeddings and, therefore, produce better accurate code models for code understanding tasks. Besides, we computed the average distance between the data points from different classes, i.e., the average distance between red points and blue points in Figure. 3. The results from Table 4 confirm that GenCode can better distinguish the two classes. Interestingly, we found that MixCode has the smallest distance result. That means even though MixCode can train code models with good performance, the code models have low confidence in their predictions, thus, the data are near the decision boundaries. This phenomenon is similar to the findings in (Verma et al. 2019), where interpolating hidden state representations yields two beneficial properties: 1) class representations are flattened into a minimal number of directions of variation, and 2) all points between these flat representations are assigned low-confidence predictions.
Robustness Evaluation
Overall robustness. Column MHM and ALERT in Table 3 summarize the attack success rate of adversarial attack methods against code models. First, we can see that similar to the observation in Accuracy Evaluation, not all data augmentation methods can improve the robustness of code models. This finding is consistent with the existing works (Dong et al. 2023b). Specifically, in the baseline methods, only RS and MixCode have a positive impact on the robustness of code models. Considering our method, GenCode consistently improves the robustness of code models and achieves the best results in 21 out of 24 cases. On average, compared to No Aug, models trained by GenCode have an 8.42% lower ASR (higher robustness). This result breaks the previous finding (Bielik and Vechev 2020) that states simply adding additional generated code to the training dataset to train code models has limited benefits in improving their robustness. On the other hand, compared to MixCode, GenCode can train code models with 4.90% higher robustness on average.
Statistical analysis. We also conduct statistical tests using T-Test and Wilcoxon signed-rank test. When comparing existing data augmentation methods, GenCode consistently exhibits a p-value $<$ 0.05 across all pre-trained models. In conclusion, GenCode can produce more adversarial robust code understanding models than other methods.
Maximum probability analysis. To explore why GenCode produces a more robust code model, we check the confidence of trained models using different code data augmentation methods on the original test data. Normally, if the DNN model is more confident in the test data, it is more difficult to attack this model using this data (Wu et al. 2018). To do so, for each dataset and code model, we collect the maximum probability of the corrected classified test data produced by the code model first. Then, we calculate the mean values of these maximum probability scores. The results of the mean values of maximum probabilities are shown in Table 5. We can see that code models trained by GenCode are more confident in the test data than models trained using other code data augmentation methods. As a result, GenCode trained code models are harder to attack.
Ablation Study
Table 6 summarizes the accuracy and robustness of trained code models using different selection strategies. We can see GenCode consistently outperforms the two variants under all testing criteria. Selecting code samples with minimum loss values is the worst way for GenCode. Interestingly, we found that GenCode (Random) performs better than all baseline methods (i.e., No Aug, Refactor, and MixCode) in terms of both accuracy and robustness improvements based on the results. This phenomenon indicates simply combining all syntax-preserving methods including code refactoring methods and syntax-breaking including text-oriented methods is already a good code data augmentation method.
Discussion

### v3

@@ 7.1 RQ1: How effective is GenCode in producing accurate code models?
Table 5: Results of statistical tests on Accuracy & Natural Robustness. A gray background highlights that the comparison is statistically significant. Statistical tests method: Wilcoxon signed-rank test. Across CodeBERT, GraphCodeBERT, CodeT5, StarCoder2, and Qwen2.5-Coder models.
DA method No Aug Refactor SR RI RS RD BT MixCode
Accuracy GenCode 1.907e-06 1.907e-06 1.907e-06 8.845e-05 1.907e-06 1.907e-06 1.907e-06 0.001
Natural Robustness GenCode 1.907e-06 - 1.907e-06 1.907e-06 1.907e-06 1.316e-04 1.907e-06 6.533e-04
Statistical analysis. We conducted statistical tests with Wilcoxon signed-rank test, across all code models. The results from the upper of Table 5 show that GenCode significantly (with p-value << 0.05) outperforms other methods including MixCode. In conclusion, GenCode is effective in training accurate code models.
Answer to RQ1: GenCode consistently outperforms baselines in training more accurate code understanding models. Specifically, for pre-trained code models, GenCode yields an average accuracy improvement of 3.06% over models trained without data augmentation and 2.92% over those trained with the SOTA MixCode method. For LLMs, GenCode achieves average improvements of 0.93% for StarCoder2 and 0.84% for Qwen2.5-Coder.
7.2 RQ2: How effective is GenCode in producing robust code models?
Adversarial robustness evaluation.
Table 6: Effectiveness of data augmentation methods w.r.t. ASR $\downarrow$ (%) on test data. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
GCJ Refactory Java250 BigCloneBench
Model DA method MHM ALERT MHM ALERT MHM ALERT MHM ALERT
No Aug 33.31 56.13 36.31 42.31 40.55 53.37 9.74 24.61
Refactor 35.56 52.84 32.64 40.89 38.35 52.56 8.55 18.29
SR 35.46 52.71 39.25 53.91 42.01 53.79 12.52 26.11
RI 40.78 65.76 38.45 45.57 43.15 55.07 8.31 30.08
RS 28.26 52.44 38.63 45.98 39.36 50.79 13.41 30.13
RD 35.65 55.46 30.23 41.32 38.31 52.42 9.58 30.15
BT 40.38 55.95 62.98 69.87 42.21 53.32 11.29 26.32
MixCode 29.28 52.53 31.04 41.01 36.14 50.67 - -
CodeBERT GenCode 22.33 46.62 23.75 31.25 29.89 40.65 7.12 15.83
No Aug 31.23 56.02 22.97 30.26 23.16 42.66 4.19 12.22
Refactor 30.31 56.58 24.26 32.57 23.13 42.42 6.05 6.17
SR 40.02 66.68 28.39 47.15 22.92 42.18 6.13 10.16
RI 43.51 65.75 25.08 32.87 22.37 42.33 2.23 6.28
RS 30.18 54.75 21.85 29.67 22.68 41.98 6.03 11.94
RD 31.73 60.92 20.35 26.98 22.39 41.84 6.18 8.27
BT 41.65 62.99 35.56 46.23 22.91 42.32 8.18 14.12
MixCode 30.26 54.49 14.89 26.53 20.03 42.01 - -
GraphCodeBERT GenCode 26.43 48.87 15.78 21.92 18.76 33.81 1.86 4.62
No Aug 30.78 55.67 22.45 29.69 31.76 49.82 10.32 22.97
Refactor 30.12 55.93 28.76 35.45 30.66 47.12 9.32 22.05
SR 39.21 51.84 28.24 36.26 28.65 46.01 15.87 25.98
RI 41.02 63.46 24.78 33.61 27.14 46.76 10.43 29.76
RS 29.35 51.74 21.32 27.76 28.45 48.45 14.33 30.43
RD 31.04 56.85 19.93 24.87 26.09 45.24 9.98 28.87
BT 39.65 54.24 33.62 43.67 29.47 45.38 10.43 25.34
MixCode 29.82 42.56 12.49 18.32 28.73 48.87 - -
CodeT5 GenCode 23.57 44.76 12.13 19.56 21.63 39.76 7.39 12.15
Column MHM and ALERT in Table 6 summarize the attack success rate of adversarial attack methods against code models. First, we can see that similar to the observation in Accuracy Evaluation, not all data augmentation methods can improve the robustness of code models. This finding is consistent with the existing works Dong et al. (2025). Specifically, in the baseline methods, only RS and MixCode have a positive impact on the robustness of code models. Considering our method, GenCode consistently improves the robustness of code models and achieves the best results in 21 out of 24 cases. On average, compared to No Aug, models trained by GenCode have an 8.42% lower ASR (higher robustness). This result breaks the previous finding Bielik and Vechev (2020) that states simply adding additional generated code to the training dataset to train code models has limited benefits in improving their robustness. On the other hand, compared to MixCode, GenCode can train code models with 4.90% higher robustness on average.
Natural robustness evaluation. Table 7 presents the result of the natural robustness evaluation. For a fair comparison, we remove Refactor from the column DA method. The results indicate that most text-oriented data augmentation methods Dong et al. (2025), such as BT and RI, do not contribute significantly to improving natural robustness. Sampling-based code augmentation methods, such as MixCode, yield limited improvements. For instance, in code multi-classification tasks including authorship attribution (GCJ), MixCode achieves an average improvement of only 0.56% compared to training without data augmentation. Although GenCode outperforms all baselines in Table 7, the improvement remains limited, particularly for LLMs. For example, GenCode achieves an average improvement of only 0.98% for StarCoder2-7B and 0.84% for Qwen2.5-Coder-7B across all tasks. This observation aligns with findings from existing studies Dong et al. (2025); Ding et al. (2024), suggesting that the effectiveness of data augmentation diminishes as model scale increases.
Table 7: Effectiveness of data augmentation methods w.r.t. test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on transformed test data. No Aug: without data augmentation. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
Model DA method GCJ Refactory Java250 BigCloneBench
No Aug 87.57 ± 0.13 92.12 ± 0.09 89.68 ± 0.16 91.61 ± 0.18
SR 62.78 ± 0.22 89.87 ± 0.15 85.51 ± 0.07 86.21 ± 0.12
RI 50.35 ± 0.19 90.23 ± 0.19 85.21 ± 0.11 92.02 ± 0.13
RS 88.23 ± 0.13 90.56 ± 0.12 89.97 ± 0.14 85.65 ± 0.11
RD 69.34 ± 0.18 93.78 ± 0.11 90.34 ± 0.13 91.87 ± 0.12
BT 75.43 ± 0.15 82.35 ± 0.21 85.23 ± 0.15 87.78 ± 0.09
MixCode 89.11 ± 0.13 95.36 ± 0.09 93.19 ± 0.12 -
CodeBERT GenCode 91.23 ± 0.14 95.63 ± 0.14 93.62 ± 0.09 92.56 ± 0.15
No Aug 87.86 ± 0.12 92.04 ± 0.03 90.58 ± 0.08 91.89 ± 0.09
SR 66.45 ± 0.09 89.12 ± 0.06 91.05 ± 0.12 89.23 ± 0.13
RI 53.87 ± 0.13 91.29 ± 0.07 90.88 ± 0.06 93.57 ± 0.15
RS 87.73 ± 0.15 92.56 ± 0.02 91.44 ± 0.11 89.02 ± 0.08
RD 73.59 ± 0.08 92.87 ± 0.05 92.01 ± 0.05 92.21 ± 0.16
BT 73.24 ± 0.11 86.05 ± 0.11 93.14 ± 0.04 90.25 ± 0.12
MixCode 88.13 ± 0.13 95.24 ± 0.12 94.05 ± 0.15 -
GraphCodeBERT GenCode 89.78 ± 0.16 95.67 ± 0.15 94.89 ± 0.08 93.89 ± 0.11
No Aug 89.31 ± 0.15 92.89 ± 0.08 91.13 ± 0.12 91.93 ± 0.13
SR 68.94 ± 0.06 90.21 ± 0.12 88.23 ± 0.08 89.43 ± 0.08
RI 58.34 ± 0.18 91.14 ± 0.14 91.13 ± 0.13 93.29 ± 0.17
RS 89.98 ± 0.19 92.89 ± 0.13 90.87 ± 0.07 88.34 ± 0.07
RD 75.57 ± 0.13 93.98 ± 0.12 91.97 ± 0.11 92.35 ± 0.14
BT 75.64 ± 0.11 88.79 ± 0.08 94.06 ± 0.16 90.87 ± 0.05
MixCode 88.45 ± 0.14 95.86 ± 0.14 93.89 ± 0.19 -
CodeT5 GenCode 91.38 ± 0.09 95.87 ± 0.11 95.06 ± 0.12 94.03 ± 0.12
No Aug 92.53 ± 0.02 94.12 ± 0.05 92.71 ± 0.11 93.21 ± 0.04
SR 82.75 ± 0.03 92.31 ± 0.07 91.87 ± 0.03 91.45 ± 0.05
RI 75.89 ± 0.05 93.19 ± 0.12 92.35 ± 0.05 93.88 ± 0.13
RS 92.13 ± 0.06 94.11 ± 0.08 91.67 ± 0.14 90.11 ± 0.06
RD 90.15 ± 0.08 94.67 ± 0.09 93.03 ± 0.02 92.89 ± 0.09
BT 85.62 ± 0.04 90.12 ± 0.13 92.55 ± 0.09 91.12 ± 0.03
MixCode 92.62 ± 0.01 95.91 ± 0.01 92.58 ± 0.06 -
StarCoder2-7B GenCode 92.91 ± 0.03 96.32 ± 0.12 93.03 ± 0.04 94.21 ± 0.12
No Aug 92.61 ± 0.03 94.35 ± 0.11 92.82 ± 0.14 93.45 ± 0.06
SR 85.67 ± 0.04 92.67 ± 0.13 91.93 ± 0.08 91.55 ± 0.08
RI 78.94 ± 0.06 93.01 ± 0.21 92.65 ± 0.06 93.75 ± 0.04
RS 91.59 ± 0.07 94.31 ± 0.09 92.04 ± 0.02 90.32 ± 0.09
RD 90.65 ± 0.11 95.01 ± 0.04 92.65 ± 0.07 93.14 ± 0.12
BT 83.65 ± 0.02 91.23 ± 0.11 92.86 ± 0.13 91.56 ± 0.05
MixCode 92.89 ± 0.09 94.89 ± 0.02 92.67 ± 0.03 -
Qwen2.5-Coder-7B GenCode 93.12 ± 0.02 96.17 ± 0.15 92.93 ± 0.09 94.38 ± 0.14
Maximum probability analysis. To explore why GenCode produces a more robust code model, we check the confidence of trained models using different code augmentation methods on the original test data. Normally, if the DNN model is more confident in the test data, it is more difficult to attack this model using this data Wu et al. (2018). To do so, for each dataset and code model, we collect the maximum probability of the corrected classified test data produced by the code model first. Then, we calculate the mean values of these maximum probability scores. The results of the mean values of maximum probabilities are shown in Table 8. We can see that code models trained by GenCode are more confident in the test data than models trained using other code augmentation methods. As a result, GenCode trained code models are harder to attack.
@@ 7.2 RQ2: How effective is GenCode in producing robust code models?
Statistical analysis. We also conduct statistical tests using Wilcoxon signed-rank test. When comparing existing data augmentation methods, GenCode consistently exhibits a p-value << 0.05 across all pre-trained code models shown from the lower of Table 5. In conclusion, GenCode can produce more adversarial robust code understanding models than other methods.
Answer to RQ2: Our statistical tests find GenCode can train more robust pre-trained code models than existing data augmentation methods. On average, code models trained using GenCode have 8.42% and 4.90% higher robustness than models trained without data augmentation and models trained using MixCode, respectively. However, for LLMs, the robustness improvement achieved by GenCode remains limited, with gains of less than 1%.
7.3 RQ3: How does the influence score affect the effectiveness of GenCode?
Table 9: Effectiveness of data augmentation methods w.r.t. test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on original test data. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
Model DA method GCJ Refactory Java250 BigCloneBench
Random 94.74 ± 0.21 98.82 ± 0.15 98.24 ± 0.11 98.27 ± 0.19
Min 93.34 ± 0.32 96.13 ± 0.21 96.43 ± 0.17 97.23 ± 0.14
CodeBERT Max 96.23 ± 0.18 99.32 ± 0.24 98.89 ± 0.09 98.93 ± 0.15
Random 95.26 ± 0.11 98.93 ± 0.18 98.68 ± 0.22 98.67 ± 0.13
Min 93.53 ± 0.17 96.31 ± 0.24 96.72 ± 0.12 97.22 ± 0.21
GraphCodeBERT Max 97.93 ± 0.14 99.45 ± 0.17 99.16 ± 0.13 99.23 ± 0.09
Random 96.24 ± 0.21 98.99 ± 0.21 98.93 ± 0.16 98.83 ± 0.08
Min 94.11 ± 0.31 98.29 ± 0.12 97.03 ± 0.23 97.01 ± 0.15
CodeT5 Max 98.85 ± 0.16 99.64 ± 0.37 99.38 ± 0.11 99.34 ± 0.12
Random 96.81 ± 0.06 99.11 ± 0.02 99.23 ± 0.04 99.31 ± 0.06
Min 95.42 ± 0.03 98.68 ± 0.04 98.47 ± 0.03 98.13 ± 0.03
StarCoder2-7B Max 98.93 ± 0.04 99.65 ± 0.02 99.41 ± 0.05 99.36 ± 0.09
Random 97.12 ± 0.05 99.12 ± 0.03 99.23 ± 0.03 99.32 ± 0.04
Min 95.86 ± 0.08 98.71 ± 0.01 98.51 ± 0.05 98.19 ± 0.07
Qwen2.5-Coder-7B Max 98.96 ± 0.04 99.65 ± 0.01 99.42 ± 0.04 99.36 ± 0.05
Table 9 and Table 10 summarize the accuracy and robustness of trained code models using different selection strategies. We can see GenCode consistently outperforms the two variants under all testing criteria. Selecting code samples with minimum loss values is the worst way for GenCode. Interestingly, we found that GenCode (Random) performs better than all baseline methods (i.e., No Aug, Refactor, and MixCode) in terms of both accuracy and robustness improvements based on the results.
Table 10: Effectiveness of data augmentation methods w.r.t. ASR $\downarrow$ (%) on test data. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
@@ 7.3 RQ3: How does the influence score affect the effectiveness of GenCode?
Model DA method MHM ALERT MHM ALERT MHM ALERT MHM ALERT
Random 25.21 50.32 27.65 36.65 32.52 44.87 8.01 18.12
Min 35.66 66.67 65.51 69.76 46.76 53.86 20.65 41.32
CodeBERT Max 22.33 42.62 22.75 29.25 29.89 40.65 7.12 15.83
Random 29.97 50.63 17.56 24.29 20.65 38.86 2.04 5.97
Min 40.55 68.23 38.73 46.98 26.87 46.67 11.45 21.23
GraphCodeBERT Max 26.43 39.87 15.78 20.92 18.76 33.81 1.86 4.62
Random 25.03 47.61 18.46 23.67 25.11 42.34 8.67 17.42
Min 38.65 59.89 35.67 48.59 32.61 49.27 18.36 25.54
CodeT5 Max 23.57 38.76 12.13 18.81 21.63 39.76 7.39 12.15
This phenomenon indicates that simply combining all syntax-preserving methods including code refactoring methods and syntax-breaking including text-oriented methods, is already a good code ugmentation method.
Answer to RQ3: Selecting code samples with maximum loss values is the best strategy among our considered variants, i.e., with minimum loss values and random selection. Besides, GenCode with random selection can outperform existing code augmentation methods.
7.4 RQ4: How does the data selection method affect the effectiveness of GenCode?
Table 11 presents the results of accuracy. Overall, LIME demonstrates limited effectiveness and generally underperforms compared to the gradient-based method (i.e.,  BADGE). Influence Function achieves better performance in scenarios involving smaller datasets, such as CodeT5 on GCJ and GraphCodeBERT on Refactory. Additionally, BADGE exhibits superior performance on LLMs, particularly in tasks such as Qwen2.5-Coder-7B on Refactory and BigCloneBench. Although GenCode outperforms other methods in 12 out of 20 cases, these results suggest that exploring combinations of existing data selection strategies could be a promising direction for future research.
Table 11: Effectiveness of data augmentation methods w.r.t. test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on original test data. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
Model Data Selection GCJ Refactory Java250 BigCloneBench
LIME 93.67 ± 0.21 98.81 ± 0.13 97.25 ± 0.19 97.45 ± 0.14
Influence Function 95.65 ± 0.19 99.38 ± 0.09 97.13 ± 0.24 97.58 ± 0.19
BADGE 96.15 ± 0.13 99.34 ± 0.22 98.21 ± 0.17 98.95 ± 0.12
CodeBERT GenCode 96.23 ± 0.18 99.32 ± 0.24 98.89 ± 0.09 98.93 ± 0.15
LIME 94.98 ± 0.07 98.12 ± 0.17 98.01 ± 0.09 97.83 ± 0.09
Influence Function 97.98 ± 0.11 99.46 ± 0.02 98.79 ± 0.08 97.79 ± 0.13
BADGE 96.88 ± 0.07 98.65 ± 0.06 99.03 ± 0.11 98.41 ± 0.04
GraphCodeBERT GenCode 97.93 ± 0.14 99.45 ± 0.17 99.16 ± 0.13 99.23 ± 0.09
LIME 96.36 ± 0.11 98.45 ± 0.06 99.03 ± 0.04 98.21 ± 0.13
Influence Function 98.97 ± 0.09 99.23 ± 0.13 99.29 ± 0.13 99.31 ± 0.12
BADGE 98.86 ± 0.18 99.54 ± 0.16 99.36 ± 0.03 99.33 ± 0.18
CodeT5 GenCode 98.85 ± 0.16 99.64 ± 0.37 99.38 ± 0.11 99.34 ± 0.12
LIME 96.96 ± 0.08 99.13 ± 0.04 99.25 ± 0.09 99.32 ± 0.04
Influence Function 97.35 ± 0.05 99.41 ± 0.06 99.31 ± 0.05 99.34 ± 0.03
BADGE 98.96 ± 0.14 99.65 ± 0.04 99.39 ± 0.12 99.35 ± 0.06
StarCoder2-7B GenCode 98.93 ± 0.04 99.65 ± 0.02 99.41 ± 0.05 99.36 ± 0.09
LIME 97.46 ± 0.05 99.25 ± 0.12 99.04 ± 0.02 99.32 ± 0.13
Influence Function 98.13 ± 0.07 99.52 ± 0.05 99.12 ± 0.07 99.35 ± 0.02
BADGE 98.98 ± 0.06 99.49 ± 0.03 99.38 ± 0.02 99.37 ± 0.03
Qwen2.5-Coder-7B GenCode 98.96 ± 0.04 99.65 ± 0.01 99.42 ± 0.04 99.36 ± 0.05
Table 12: Effectiveness of data augmentation methods w.r.t. test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on transformed test data. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
Model Data Selection GCJ Refactory Java250 BigCloneBench
LIME 89.13 ± 0.07 95.41 ± 0.11 93.21 ± 0.14 91.68 ± 0.12
Influence Function 90.34 ± 0.05 95.51 ± 0.07 93.35 ± 0.12 92.17 ± 0.15
BADGE 91.31 ± 0.09 95.61 ± 0.05 93.61 ± 0.03 92.64 ± 0.11
CodeBERT GenCode 91.23 ± 0.14 95.63 ± 0.14 93.62 ± 0.09 92.56 ± 0.15
LIME 88.24 ± 0.06 95.35 ± 0.07 94.24 ± 0.02 93.62 ± 0.08
Influence Function 88.97 ± 0.08 95.62 ± 0.03 94.32 ± 0.07 93.58 ± 0.12
BADGE 90.01 ± 0.04 95.88 ± 0.05 94.87 ± 0.11 93.96 ± 0.07
GraphCodeBERT GenCode 89.78 ± 0.16 95.67 ± 0.15 94.89 ± 0.08 93.89 ± 0.11
LIME 90.45 ± 0.13 95.85 ± 0.13 94.32 ± 0.04 93.43 ± 0.05
Influence Function 91.04 ± 0.16 95.94 ± 0.11 94.68 ± 0.09 93.65 ± 0.09
BADGE 91.35 ± 0.14 96.14 ± 0.15 95.14 ± 0.09 93.95 ± 0.14
CodeT5 GenCode 91.38 ± 0.09 95.87 ± 0.11 95.06 ± 0.12 94.03 ± 0.12
LIME 92.72 ± 0.16 94.56 ± 0.14 92.67 ± 0.12 93.91 ± 0.09
Influence Function 92.89 ± 0.03 95.13 ± 0.19 92.86 ± 0.11 94.12 ± 0.15
BADGE 92.93 ± 0.04 96.24 ± 0.17 93.12 ± 0.13 94.19 ± 0.19
StarCoder2-7B GenCode 92.91 ± 0.03 96.32 ± 0.12 93.03 ± 0.04 94.21 ± 0.12
LIME 92.94 ± 0.07 95.13 ± 0.04 92.68 ± 0.12 93.87 ± 0.05
Influence Function 92.88 ± 0.19 95.78 ± 0.09 92.76 ± 0.08 94.19 ± 0.06
BADGE 93.09 ± 0.05 96.19 ± 0.02 92.97 ± 0.11 94.32 ± 0.05
Qwen2.5-Coder-7B GenCode 93.12 ± 0.02 96.17 ± 0.15 92.93 ± 0.09 94.38 ± 0.14
An interesting observation from the natural robustness results in Table 12 is that, in most cases, the gradient-based method achieves better robustness improvements than GenCode. Specifically, BADGE outperforms GenCode by integrating gradient-based uncertainty with diversity via K-means clustering, thereby facilitating the selection of informative and representative samples Ash et al. (2019).
Answer to RQ4: GenCode achieves better improvements in accuracy compared to other data selection methods. However, the gradient-based method demonstrates superior performance in enhancing robustness. Investigating the trade-offs between GenCode and gradient-based methods presents an interesting direction for future research.
8 Discussion
8.1 Computational cost of augmented data
We conduct a case study to evaluate the computational cost of model training with different data augmentation methods. The results are visualized in Fig. 6. We observe that GenCode incurs the highest computational cost compared to other data augmentation methods including code refactoring, text-oriented, and MixCode. Interestingly, text-oriented data augmentation methods consistently exhibit lower training-time cost than code refactoring methods. In particular, Random Deletion (RD) achieves the best trade-off between performance and efficiency.
Although Back Translation (BT) performs relatively well across some tasks (e.g., authorship attribution and bug detection), it requires significant preprocessing time to generate the augmented data before model training, making it less practical. Therefore, we do not recommend BT as a preferred augmentation strategy in time-sensitive scenarios.
(a) Authorship attribution
(b) Bug detection
(c) Problem classification
(d) Clone detection
Figure 6: Computational cost of CodeBERT using different code augmentation methods in each task.
In conclusion, rule-based augmentation methods (e.g., code refactoring or text-oriented data augmentation methods) typically require less computation during model training, but may incur the preprocessing overhead. In contrast, sample-based methods such as MixCode do not require data preparation in advance and can be flexibly applied during training. While GenCode consistently achieves the best accuracy and robustness, it comes with the highest computational cost. Exploring the trade-off between performance and efficiency is an important direction for future work.
8.2 On the Data Distribution of Augmented Data
One common concern in data augmentation is the risk of introducing samples that deviate from the original data distribution, potentially harming model generalization. While augmented data may not strictly follow the original distribution, prior studies (e.g., Mixup Zhang et al. (2018), CutMix Yun et al. (2019)) have shown that moderate, controlled deviations can actually enhance generalization. To mitigate distributional shift, GenCode follows a principled two-stage pipeline: (1) controlled data generation, and (2) loss-based importance-guided selection. In the first stage, GenCode applies a combination of semantic-preserving code refactorings (e.g., variable renaming, field enhancement) and lightweight text-oriented data augmentation methods(e.g., random swap, synonym replacement). These transformations are designed to preserve or minimally alter program semantics, ensuring that the generated data remains close to the original distribution.
In the second stage, GenCode evaluates each augmented sample based on its training loss under the current model. Only samples with the highest loss, those that are potentially more informative, are selected for training, while others are discarded. This loss-based filtering mechanism promotes alignment with the model’s effective learning region and helps avoid training on overly noisy or unrepresentative data.
8.3 Limitations